a stick that is 3 feet long is divided into two uneven pieces so that the ratio of the length of the short piece to the length of the long piece is equal to the ratio of the length of the long piece to the length of the original stick. How long is the length of the short piece?
my work:
a=short stick
b=long stick
system of equations: a+b=3 feet(original stick)36in
a/b=b/3

Answer by richwmiller(17219)

Using your equations
a+b=3
a/b=b/3
a=1.1459, b=1.8541